OpenStudy (anonymous):

Solve the following systems of equations using the graphing method. What type of system is it? Name the solution if there is one. 2x-2y=8 x-y=4

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

2x-2y=8 +2y +2y 2x=2y+8 /2 /2 x=y+4 x-y=4 +x +x y=x+4

OpenStudy (anonymous):

@megan_tanuis

OpenStudy (rob1525):

graph and find the intersection

OpenStudy (anonymous):

I'm confused on how to graph. is my work above correct

OpenStudy (anonymous):

You subtract 2x, not add. Whatever the sign is, you do the opposite so it cancels out

OpenStudy (rob1525):

I think their both reduce to y=x-4

OpenStudy (rob1525):

no intersection means no solution

OpenStudy (aum):

2x-2y=8 divide by 2: x - y = 4 add y: x = y + 4 subtract 4: x - 4 = y or y = x - 4 x-y=4 add y: x = y + 4 subtract 4: x - 4 = y or y = x - 4 They are both one and the same equation. You will get just one line. Any point on the line will satisfy both the equations. You have infinte number of solutions.

OpenStudy (anonymous):

2x-2y=8 -2x -2x -2y=-2x+8 /-2 /-2 y=x-4

OpenStudy (anonymous):

i see my mistake

OpenStudy (anonymous):

There, now you have it

OpenStudy (anonymous):

Do you still need help with graphing?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

You get the same equation. So it's just one line. Plot your first point at -4 since that is your y-intercept. Then from that point, you go up one, over one.

OpenStudy (anonymous):



OpenStudy (anonymous):

@megan_tanuis

OpenStudy (anonymous):

What is the second circle for?

OpenStudy (anonymous):

you said up one over to right?

OpenStudy (anonymous):

Like this

OpenStudy (anonymous):

what do you use to make your graphs? lol they look much better than mine

OpenStudy (anonymous):

But yes, I can see it now. You continue up the line with that.

OpenStudy (anonymous):

and it is infinite?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If you need anymore help, I will be on tomorrow after 11am

OpenStudy (anonymous):

okay i do have some more is that 11am eastern?

OpenStudy (anonymous):

Yeah